# Fundamental limitations on the recoverability of quantum processes

### v1

@@ II Preliminaries: Standard notations, definitions, and facts
where $\mathsf{X}^{t}$ denote the transpose of operator $\mathsf{X}$, $\mathcal{N}\in\mathcal{L}(A,B)$, and $\widehat{\mathsf{C}}_{\mathcal{N}}$ is the Choi operator for $\mathcal{N}$. Before proceeding further we first find the adjoints of the maps $\mathsf{T}$ and $\mathsf{T}^{{}^{\prime}}$.
@@
The adjoint $\mathsf{T}^{*}:\mathcal{L}(C\otimes D)\rightarrow\mathcal{L}(A\otimes B)$ of the representing map $\mathsf{T}$ of a supermap $\Theta:\mathcal{L}(A,B)\rightarrow\mathcal{L}(C,D)$, is given by
@@
This result, which tells that the identification $\mathsf{T}\leftrightarrow\Theta$ is a ${}^{\ast}$-isomorphism, was proved in Gou19a . We provide an alternate proof of this result in Appendix B.4.
@@
The adjoint $\mathsf{T}^{{}^{\prime}*}:\mathcal{L}\left({C}\otimes{D}\right)\rightarrow\mathcal{L}\left({A}\otimes{B}\right)$ of the map $\mathsf{T}^{{}^{\prime}}$, which is defined via Eq. (20), is given by
@@
The adjoint map $\mathfrak{T}^{*}$.– We already know that $\mathsf{T}^{*}\left(\widehat{\mathsf{C}}_{\mathcal{M}}\right)=\widehat{\mathsf{C}}_{\mathrm{\Theta}^{*}(\mathcal{M})}$. As the map $\mathsf{T}\leftrightarrow\mathrm{\Theta}$ is product preserving, we have $T^{-1}\left(\widehat{\mathsf{C}}_{\mathcal{M}}\right)=\widehat{\mathsf{C}}_{\mathrm{\Theta}^{-1}(\mathcal{M})}$. From this relation we derive the action of $\mathfrak{T}^{*}:\mathcal{L}({C}\otimes{D})\rightarrow\mathcal{L}({A}\otimes{B})$ on $\widehat{\mathsf{C}}_{\mathcal{M}}^{\Phi}$ as
@@
Representing map for super-supermaps.– Let us denote by $\mathring{\mathbf{\Lambda}}:\mathscr{L}\left[\mathcal{L}\left(A,B\right),\mathcal{L}\left(C,D\right)\right]\rightarrow\mathscr{L}\left[\mathcal{L}\left(A^{\prime},B^{\prime}\right),\mathcal{L}\left(C^{\prime},D^{\prime}\right)\right]$ a higher order process transforming supermaps to supermaps. We call it a super-supermap. Now, we can define its representing supermap $\mathbb{T}_{\mathring{\mathbf{\Lambda}}}:\mathcal{L}\left(A,B\right)\otimes\mathcal{L}\left(C,D\right)\rightarrow\mathcal{L}\left(A^{\prime},B^{\prime}\right)\otimes\mathcal{L}\left(C^{\prime},D^{\prime}\right)$ by
@@
where $\mkern 2.0mu\overline{\mkern-2.0mu\mathcal{N}\mkern-2.0mu}\mkern 2.0mu$ represents the complex conjugate of the map $\mathcal{N}$ defined as $\mkern 2.0mu\overline{\mkern-2.0mu\mathcal{N}\mkern-2.0mu}\mkern 2.0mu:=\sum_{ijkl}\bar{\alpha}_{ijkl}\mathcal{E}^{A^{\prime}\rightarrow B^{\prime}}_{ijkl}$ when $\mathcal{N}$ is written as $\mathcal{N}:=\sum_{ijkl}\alpha_{ijkl}\mathcal{E}^{A^{\prime}\rightarrow B^{\prime}}_{ijkl}$ and ${}^{\ast}$ represents the adjoint as defined in Eq. (7). The trace of a map $\mathcal{N}:\mathcal{L}(A^{\prime})\rightarrow\mathcal{L}(A^{\prime})$ is defined by $\operatorname{tr}[\mathcal{N}]=\sum_{ij}\left<e^{A^{\prime}}_{ij},\mathcal{N}\left(e^{A^{\prime}}_{ij}\right)\right>$ with $e^{A^{\prime}}_{ij}$ being the canonical basis of the space $A^{\prime}$. Using the form of Eq. (13) for $\mathbb{T}_{\mathring{\mathbf{\Lambda}}}$ in Eq. (34), we expect a representation similar to Eq. (13) to hold for a super-superchannel $\mathring{\mathbf{\Lambda}}$.
@@ III Entropy change of quantum channels
Distinguishing two different quantum states is one of the most basic tasks in the quantum information theory as multitude of problems in the quantum information theory can be recast as state distinguishability problems Hel (76). In the asymptotic limit of quantum state discrimination the minimum error in distinguishing two quantum states is expressed neatly by quantum Chernoff bound ${}^{+}$ (07), which is very closely related to the relative entropy between the states Ved (02). This provides an operational meaning to the relative entropy between two quantum states. On the other hand, relative entropy plays a pivotal role in proving a plethora of results such as the strong subaddivity of the von Neumann entropy NC (00) and recoverability of quantum states affected by quantum noise OP (93). In fact, the quantum relative entropy can be treated as a parent quantity for analysis as many other information theoretic quantities such as the von Neumann entropy, the quantum mutual information NC (00), and the free energy ${}^{+}$ (13) can be re-expressed in terms of relative entropy. Our interest here is in the von Neumann entropy. The von Neumann entropy is a measure of randomness or uncertainty present in a quantum system and has a variety of uses in quantum information theory, e.g., it is crucial in the theory of entanglement Sch (96), quantum communications Hol (98); YO (93), and quantum cryptography CBTW (17). It is a natural generalization of Shannon entropy Sha (48). We can write von Neumann entropy in terms of the quantum relative entropy as follows
@@
Recall that under a positive trace preserving sub-unital map $\mathcal{N}$, the entropy of a quantum state never decreases i.e., $S(\mathcal{N}(\rho))\geq S(\rho)$. We want to investigate if a similar statement holds for the entropy of quantum channels as well. We note that the generalization of the identity matrix is the completely depolarizing map (see Definition 2) and hence the generalization of sub-unitality for a supermap $\mathrm{\Theta}$ is captured by $\mathrm{\Theta}(\mathcal{R})\leq\mathcal{R}$, where the ordering between maps is defined in Proposition 7. Here, we ask whether the entropy of a quantum channel is nondecreasing under those quantum superchannels $\mathrm{\Theta}$ for which $\mathrm{\Theta}(\mathcal{R}_{A\rightarrow B})\leq\mathcal{R}_{C\rightarrow D}$ holds. To answer this question, first we rewrite the definition of entropy of a quantum channel which is in a similar spirit to Eq. (35) (see Definition 5 ), but to do that we first need the notion of relative entropy between a quantum channel and a completely positive (CP) map. This is because, in the settings of quantum channels, the generalization of identity is the completely depolarising map. In Proposition 8 we have shown that if $\Theta(\mathcal{R}_{A\rightarrow B})\leq\mathcal{R}_{C\rightarrow D}$, then the entropy of channels is non-decreasing under the action of this superchannel, i.e., $S\left[\Theta(\mathcal{N}_{A\rightarrow B})\right]\geq S\left[\mathcal{N}_{A\rightarrow B}\right]$. We also observe that a necessary condition for the relation $S\left[\Theta(\mathcal{N}_{A\rightarrow B})\right]\geq S\left[\mathcal{N}_{A\rightarrow B}\right]$ to hold is $|B|\leq|D|$.
@@ III.1 Data processing inequality
if the pure states that realize the supremum in the relative entropy functional $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ and $D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]$ have the full rank marginals.
@@
Consider a quantum channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ and two completely positive maps $\mathcal{M}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ and ~$\widetilde{\mathcal{M}}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ with ordering $-~\widetilde{\mathcal{M}}\geq\mathcal{M}$, then we have
@@ III.2 Entropy of quantum channels and its properties
where the above ordering means that the linear map $\mathcal{R}_{C\rightarrow D}-\Theta(\mathcal{R}_{A\rightarrow B})$ is completely positive.
@@
Let $\mathcal{N}_{A\rightarrow B}$ be a replacer channel, i.e., for all $\rho\in\mathcal{D}(A)$, it satisfies $\mathcal{N}_{A\rightarrow B}(\rho)=\sigma_{0}$ for a fixed $\sigma_{0}\in\mathcal{D}(B)$. Then $S(\mathcal{N}_{A\rightarrow B})=0$ when $\sigma_{0}$ is a pure state and $S(\mathcal{N}_{A\rightarrow B})$ is maximum i.e., $S(\mathcal{N}_{A\rightarrow B})=\log|B|$ when $\sigma_{0}=\frac{1}{|B|}\mathbbm{1}_{B}$, i.e., when $\mathcal{N}_{A\rightarrow B}$ is completely depolarizing channel.
@@
Another crucial property that the entropy functional of a quantum channel has to satisfy is that it should be nondecreasing under the action of a random unitary superchannel. The definition used in Gou19b satisfies this property as a corollary of the fact that the channel entropy always increases under those quantum superchannel which preserves the completely depolarizing channel. Here we generalize this result and show that the entropy functional is nondecreasing under completely depolarizing channel subpreserving superchannels. Random isometry superchannels are one of the examples of such class of superchannels and we prove this as the following proposition.
@@
The Proposition 8 is an expression for the entropy gain of quantum channels under the action of $\mathcal{R}$-subpreserving superchannels. As a special case, this proposition yields an expression for entropy gain of quantum channels under the action of random isometry superchannels, generalizing the result of entropy gain of quantum states under noisy operations. The proof relies on the fact that the random isometry superchannels are sub-completely randomizing channel-preserving superchannels (see Proposition 7). Although it is clear that the entropy of quantum channels is nondecreasing under the action of random isometry superchannels it does not provide any estimate of change of entropy in terms of some remainder terms. In the following, we study the entropy gain of quantum channels in detail and estimate a remainder term for entropy gain for a class of quantum channels where we assume that the state that realizes the supremum has full rank reduced state. Note that the tele-covariant channels are an example of such class of channels where the supremum occurs at a maximally entangled state for which the reduced statte is proportional to the identity operator. More specifically, we have the following theorem, which is one of our main results.
@@
Observe that $\Delta^{\prime}\left[|A|^{-1}\Psi^{+}_{\widetilde{A}},|C|^{-1}\Psi^{+}_{\widetilde{C}}\right]$ is already a constant in this case and equal to $\log\widetilde{d}$, where $\widetilde{d}=|A|/|C|$. Further, we have $\left(\mathsf{C}_{\mathcal{N}}\right)_{\alpha}=\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathsf{C}_{\mathrm{\Theta}^{*}\circ\mathrm{\Theta}(\mathcal{N})}\right)^{\alpha}$, where $\alpha=\norm{\mathsf{C}_{\Theta^{*}(\mathcal{R})}}_{\infty}$. Notice that if superchannel $\Theta$ is such that $|A|\geq|C|$, then we always have $\log\widetilde{d}\geq 0$. Note that $D\left(\mathsf{C}_{\mathcal{N}}\;\middle\|\;\left(\mathsf{C}_{\mathcal{N}}\right)_{\alpha}\right)\geq 0$ imposes the condition that $\operatorname{tr}\left(\left(\mathsf{C}_{\mathcal{N}}\right)_{\alpha}\right)\leq 1$, which one can arrive using data processing inequality. Thus, we can conclude that all tele-covariant channel-preserving superchannels satisfying these two conditions must be channel entropy nondecreasing.
@@ IV Sufficiency for the recoverability of quantum channels
where $\mathcal{N}^{*}$ is the adjoint map of $\mathcal{N}$. In recent years, some refinement of data processing inequality have also been proposed in terms of a recovery map Wil (15); ${}^{+}$ (18); BLW (15); STH (16). In  ${}^{+}$ (18), it was shown that there exists a universal recovery map (channel) $\mathcal{P}^{\mathrm{R}}$ such that the monotonicity of the quantum relative entropy between two arbitrary quantum states $\rho$ and $\sigma$ under the action of an arbitrary quantum channel $\mathcal{N}\in\mathcal{L}(A,B)$ can be refined as
@@
where $F(\rho,\sigma):=\norm{\sqrt{\rho}\sqrt{\sigma}}^{2}_{1}$ is the fidelity between $\rho,\sigma\in\mathcal{D}(A)$. Here, the recovery channel $\mathcal{P}^{\mathrm{R}}$ is given by ${}^{+}$ (18) (see also Wil (13))
@@
where $\xi\in\mathcal{D}\left(A\right)$, $\mathsf{X}\in\mathcal{L}(B)$, $\Pi_{\mathcal{N}(\sigma)}$ is the projection onto the support of $\mathcal{N}(\sigma)$, and $\alpha_{t}\equiv\pi\left(\cosh(\pi t)+1\right)^{-1}/2$ can be interpreted as probability density function on $t\in\mathbb{R}$. And $\mathcal{P}^{t}_{\sigma,\mathcal{N}}\equiv\mathcal{U}_{\sigma,-t}\circ\mathcal{P}_{\sigma,\mathcal{N}}\circ\mathcal{U}_{\mathcal{N}(\sigma),t}$, where $\mathcal{U}_{\sigma,t}$ is a partial isometric map defined as $\mathcal{U}_{\sigma,t}(\mathsf{X})=\sigma^{it}\mathsf{X}\sigma^{-it}$. The fidelity between any two arbitrary states $\rho,\sigma$ is always bounded as $F(\rho,\sigma)\in[0,1]$, with $F(\rho,\sigma)=1$ iff $\rho=\sigma$ and $F(\rho,\sigma)=0$ iff $\rho\perp\sigma$. The term $\operatorname{tr}\left(\left(\mathbbm{1}_{A}-\Pi_{\mathcal{N}(\sigma)}\right)\mathsf{X}\right)\xi$ present in the above equation is there to ensure $\mathcal{P}^{\mathrm{R}}$ is trace preserving. It can be easily checked that the recovery channel $\mathcal{P}^{\mathrm{R}}$ perfectly recovers $\sigma$ from $\mathcal{N}(\sigma)$, i.e., $\left(\mathcal{P}^{\mathrm{R}}\circ\mathcal{N}\right)(\sigma)=\sigma$, and also if $\rho$ and $(\mathcal{P}^{\mathrm{R}}\circ\mathcal{N})(\rho)$ are close then $-\log F\left(\rho,(\mathcal{P}^{\mathrm{R}}\circ\mathcal{N})(\rho)\right)$ will be a small number. So, the above refinement says that it is possible to perform a recovery operation in which one can perfectly recover one state and approximately recover the other states if the change in the quantum relative entropy between the two quantum states before and after the action of the quantum channel is small Wil (15); ${}^{+}$ (18).
@@
Let $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a superchannel and $\mathcal{N}_{A\rightarrow A^{\prime}}$, and $\mathcal{M}_{A\rightarrow A^{\prime}}$ be two quantum channels. Let $\Psi_{\widetilde{A}A}\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$ and $\Phi_{\widetilde{C}C}\in\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$ be the pure states that realize the supremum in the relative entropy functional $D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]$ and $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$, respectively. Then we have

### v2

@@ II Preliminaries: Standard notations, definitions, and facts
where $\mathsf{X}^{t}$ denote the transpose of operator $\mathsf{X}$, $\mathcal{N}\in\mathcal{L}(A,B)$, and $\widehat{\mathsf{C}}_{\mathcal{N}}$ is the Choi operator for $\mathcal{N}$. Before proceeding further we first find the adjoints of the maps $\mathsf{T}$ and $\mathsf{T}{{}^{\prime}}$.
@@
The adjoint $\mathsf{T}^{*}:\mathcal{L}(C\otimes D)\rightarrow\mathcal{L}(A\otimes B)$ of the representing map $\mathsf{T}$ of a supermap $\Theta:\mathcal{L}(A,B)\rightarrow\mathcal{L}(C,D)$ is given by
@@
This result, which tells that the identification $\mathsf{T}\leftrightarrow\Theta$ is a -isomorphism, was proved in Gou19a . We provide an alternate proof of this result in Appendix B.5.
@@
The adjoint $\mathsf{T}^{{}^{\prime}*}:\mathcal{L}\left({C}\otimes{D}\right)\rightarrow\mathcal{L}\left({A}\otimes{B}\right)$ of the map $\mathsf{T}{{}^{\prime}}$, which is defined via Eq. (20), is given by
@@
The adjoint map $\mathfrak{T}^{*}$.— We already know that $\mathsf{T}^{*}\left(\widehat{\mathsf{C}}_{\mathcal{M}}\right)=\widehat{\mathsf{C}}_{\mathrm{\Theta}^{*}(\mathcal{M})}$. As the map $\mathsf{T}\leftrightarrow\mathrm{\Theta}$ is product preserving, we have $T^{-1}\left(\widehat{\mathsf{C}}_{\mathcal{M}}\right)=\widehat{\mathsf{C}}_{\mathrm{\Theta}^{-1}(\mathcal{M})}$. From this relation we derive the action of $\mathfrak{T}^{*}:\mathcal{L}({C}\otimes{D})\rightarrow\mathcal{L}({A}\otimes{B})$ on $\widehat{\mathsf{C}}_{\mathcal{M}}^{\Phi}$ as
@@
Representing map for super-supermaps.— Let us denote by $\mathring{\mathbf{\Lambda}}:\mathscr{L}\left[\mathcal{L}\left(A,B\right),\mathcal{L}\left(C,D\right)\right]\rightarrow\mathscr{L}\left[\mathcal{L}\left(A^{\prime},B^{\prime}\right),\mathcal{L}\left(C^{\prime},D^{\prime}\right)\right]$ a higher order process transforming supermaps to supermaps. We call it a super-supermap. Now, we can define its representing supermap $\mathbb{T}_{\mathring{\mathbf{\Lambda}}}:\mathcal{L}\left(A,B\right)\otimes\mathcal{L}\left(C,D\right)\rightarrow\mathcal{L}\left(A^{\prime},B^{\prime}\right)\otimes\mathcal{L}\left(C^{\prime},D^{\prime}\right)$ by
@@
where $\mkern 2.0mu\overline{\mkern-2.0mu\mathcal{N}\mkern-2.0mu}\mkern 2.0mu$ represents the complex conjugate of the map $\mathcal{N}$ defined as $\mkern 2.0mu\overline{\mkern-2.0mu\mathcal{N}\mkern-2.0mu}\mkern 2.0mu:=\sum_{ijkl}\bar{\alpha}_{ijkl}\mathcal{E}^{A^{\prime}\rightarrow B^{\prime}}_{ijkl}$ when $\mathcal{N}$ is written as $\mathcal{N}:=\sum_{ijkl}\alpha_{ijkl}\mathcal{E}^{A^{\prime}\rightarrow B^{\prime}}_{ijkl}$ and represents the adjoint as defined in Eq. (7). The trace of a map $\mathcal{N}:\mathcal{L}(A^{\prime})\rightarrow\mathcal{L}(A^{\prime})$ is defined by $\operatorname{tr}[\mathcal{N}]=\sum_{ij}\left<e^{A^{\prime}}_{ij},\mathcal{N}\left(e^{A^{\prime}}_{ij}\right)\right>$ with $e^{A^{\prime}}_{ij}$ being the canonical basis of the space $A^{\prime}$. Using the form of Eq. (13) for $\mathbb{T}_{\mathring{\mathbf{\Lambda}}}$ in Eq. (34), we expect a representation similar to Eq. (13) to hold for a super-superchannel $\mathring{\mathbf{\Lambda}}$.
@@ III Entropy change of quantum channels
Distinguishing two different quantum states is one of the most basic tasks in the quantum information theory as multitude of problems in the quantum information theory can be recast as state distinguishability problems Hel (76). In the asymptotic limit of quantum state discrimination the minimum error in distinguishing two quantum states is expressed neatly by quantum Chernoff bound ACMnT+ (07), which is very closely related to the relative entropy between the states Ved (02). This provides an operational meaning to the relative entropy between two quantum states. On the other hand, relative entropy plays a pivotal role in proving a plethora of results such as the strong subaddivity of the von Neumann entropy NC (00) and recoverability of quantum states affected by quantum noise OP (93). In fact, the quantum relative entropy can be treated as a parent quantity for analysis as many other information theoretic quantities such as the von Neumann entropy, the quantum mutual information NC (00), and the free energy BaHO+ (13) can be re-expressed in terms of relative entropy. Our interest here is in the von Neumann entropy. The von Neumann entropy is a measure of randomness or uncertainty present in a quantum system and has a variety of uses in quantum information theory, e.g., it is crucial in the theory of entanglement Sch (96), quantum communications Hol (98); YO (93), and quantum cryptography CBTW (17). It is a natural generalization of Shannon entropy Sha (48). We can write von Neumann entropy in terms of the quantum relative entropy as follows
@@
Recall that under a positive trace preserving sub-unital map $\mathcal{N}$, the entropy of a quantum state never decreases i.e., $S(\mathcal{N}(\rho))\geq S(\rho)$. We want to investigate if a similar statement holds for the entropy of quantum channels as well. We note that the generalization of the identity matrix is the completely depolarizing map (see Definition 2) and hence the generalization of sub-unitality for a supermap $\mathrm{\Theta}$ is captured by $\mathrm{\Theta}(\mathcal{R})\leq\mathcal{R}$, where the ordering between maps is defined in Proposition 7. Here, we ask whether the entropy of a quantum channel is nondecreasing under those quantum superchannels $\mathrm{\Theta}$ for which $\mathrm{\Theta}(\mathcal{R}_{A\rightarrow B})\leq\mathcal{R}_{C\rightarrow D}$ holds. To answer this question, first we rewrite the definition of entropy of a quantum channel which is in a similar spirit to Eq. (35) (see Definition 5 ), but to do that we first need the notion of relative entropy between a quantum channel and a completely positive (CP) map. This is because, in the settings of quantum channels, the generalization of identity is the completely depolarising map. In Proposition 8, we have shown that if $\Theta(\mathcal{R}_{A\rightarrow B})\leq\mathcal{R}_{C\rightarrow D}$, then the entropy of channels is non-decreasing under the action of this superchannel, i.e., $S\left[\Theta(\mathcal{N}_{A\rightarrow B})\right]\geq S\left[\mathcal{N}_{A\rightarrow B}\right]$. We also observe that a necessary condition for the relation $S\left[\Theta(\mathcal{N}_{A\rightarrow B})\right]\geq S\left[\mathcal{N}_{A\rightarrow B}\right]$ to hold is $|B|\leq|D|$.
@@ III.1 Data processing inequality
if the pure states that realize the supremum in the relative entropy functional $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ and $D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]$ have the full rank marginals, and the representing map of $\Theta$ is a trace non increasing map.
@@
Consider a quantum channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ and two completely positive maps $\mathcal{M}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ and ~$\widetilde{\mathcal{M}}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ with ordering $\widetilde{\mathcal{M}}\geq\mathcal{M}$, then we have
@@ III.2 Entropy of quantum channels and its properties
where the above ordering means that the map $\mathcal{R}_{C\rightarrow D}-\Theta(\mathcal{R}_{A\rightarrow B})$ is completely positive.
@@
Let $\mathcal{N}_{A\rightarrow B}$ be a replacer channel, i.e., for all $\rho\in\mathcal{D}(A)$, it satisfies $\mathcal{N}_{A\rightarrow B}(\rho)=\sigma_{0}$ for a fixed $\sigma_{0}\in\mathcal{D}(B)$. Then $S(\mathcal{N}_{A\rightarrow B})=0$ when $\sigma_{0}$ is a pure state and $S(\mathcal{N}_{A\rightarrow B})$ is maximum i.e., $S(\mathcal{N}_{A\rightarrow B})=\log|B|$ when $\sigma_{0}=\frac{1}{|B|}\mathbbm{1}_{B}$, i.e., when $\mathcal{N}_{A\rightarrow B}$ is the completely depolarizing channel.
@@
Another crucial property that the entropy functional of a quantum channel has to satisfy is that it should be nondecreasing under the action of a random unitary superchannel. The definition used in Gou19b satisfies this property as a corollary of the fact that the channel entropy always increases under those quantum superchannel which preserves the completely depolarizing channel. Here we generalize this result and show that the entropy functional is nondecreasing under completely depolarizing map subpreserving superchannels. Random isometry superchannels are one of the examples of such class of superchannels and we prove this as the following proposition.
@@
The Proposition 8 is an expression for the entropy gain of quantum channels under the action of $\mathcal{R}$-subpreserving superchannels. As a special case, this proposition yields an expression for entropy gain of quantum channels under the action of random isometry superchannels, generalizing the result of entropy gain of quantum states under noisy operations. The proof relies on the fact that the random isometry superchannels are sub-completely randomizing map-preserving superchannels (see Proposition 7). Although it is clear that the entropy of quantum channels is nondecreasing under the action of random isometry superchannels it does not provide any estimate of change of entropy in terms of some remainder terms. In the following, we study the entropy gain of quantum channels in detail and estimate a remainder term for entropy gain for a class of quantum channels where we assume that the state that realizes the supremum has full rank reduced state. Note that the tele-covariant channels are an example of such class of channels where the supremum occurs at a maximally entangled state for which the reduced state is proportional to the identity operator. More specifically, we have the following theorem, which is one of our main results.
@@
Observe that $\Delta^{\prime}\left[|A|^{-1}\Psi^{+}_{\widetilde{A}},|C|^{-1}\Psi^{+}_{\widetilde{C}}\right]$ is already a constant in this case and equal to $\log\widetilde{d}$, where $\widetilde{d}=|A|/|C|$. Further, we have $\left(\mathsf{C}_{\mathcal{N}}\right)_{\alpha}=\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathsf{C}_{\mathrm{\Theta}^{*}\circ\mathrm{\Theta}(\mathcal{N})}\right)^{\alpha}$, where $\alpha=\norm{\mathsf{C}_{\Theta^{*}(\mathcal{R})}}_{\infty}$. Notice that if superchannel $\Theta$ is such that $|A|\geq|C|$, then we always have $\log\widetilde{d}\geq 0$. Note that if $\operatorname{tr}\left(\left(\mathsf{C}_{\mathcal{N}}\right)_{\alpha}\right)\leq 1$, then we have $D\left(\mathsf{C}_{\mathcal{N}}\;\middle\|\;\left(\mathsf{C}_{\mathcal{N}}\right)_{\alpha}\right)\geq 0$.
@@ IV Sufficiency for the recoverability of quantum channels
where $\mathcal{N}^{*}$ is the adjoint map of $\mathcal{N}$. In recent years, some refinement of data processing inequality have also been proposed in terms of a recovery map Wil (15); JRS+ (18); BLW (15); STH (16). In  JRS+ (18), it was shown that there exists a universal recovery map (channel) $\mathcal{P}^{\mathrm{R}}$ such that the monotonicity of the quantum relative entropy between two arbitrary quantum states $\rho$ and $\sigma$ under the action of an arbitrary quantum channel $\mathcal{N}\in\mathcal{L}(A,B)$ can be refined as
@@
where $F(\rho,\sigma):=\norm{\sqrt{\rho}\sqrt{\sigma}}^{2}_{1}$ is the fidelity between $\rho,\sigma\in\mathcal{D}(A)$. Here, the recovery channel $\mathcal{P}^{\mathrm{R}}$ is given by JRS+ (18) (see also Wil (13))
@@
where $\xi\in\mathcal{D}\left(A\right)$, $\mathsf{X}\in\mathcal{L}(B)$, $\Pi_{\mathcal{N}(\sigma)}$ is the projection onto the support of $\mathcal{N}(\sigma)$, and $\alpha_{t}\equiv\pi\left(\cosh(\pi t)+1\right)^{-1}/2$ can be interpreted as probability density function on $t\in\mathbb{R}$. And $\mathcal{P}^{t}_{\sigma,\mathcal{N}}\equiv\mathcal{U}_{\sigma,-t}\circ\mathcal{P}_{\sigma,\mathcal{N}}\circ\mathcal{U}_{\mathcal{N}(\sigma),t}$, where $\mathcal{U}_{\sigma,t}$ is a partial isometric map defined as $\mathcal{U}_{\sigma,t}(\mathsf{X})=\sigma^{it}\mathsf{X}\sigma^{-it}$. The fidelity between any two arbitrary states $\rho,\sigma$ is always bounded as $F(\rho,\sigma)\in[0,1]$, with $F(\rho,\sigma)=1$ iff $\rho=\sigma$ and $F(\rho,\sigma)=0$ iff $\rho\perp\sigma$. The term $\operatorname{tr}\left(\left(\mathbbm{1}_{A}-\Pi_{\mathcal{N}(\sigma)}\right)\mathsf{X}\right)\xi$ present in the above equation is there to ensure $\mathcal{P}^{\mathrm{R}}$ is trace preserving. It can be easily checked that the recovery channel $\mathcal{P}^{\mathrm{R}}$ perfectly recovers $\sigma$ from $\mathcal{N}(\sigma)$, i.e., $\left(\mathcal{P}^{\mathrm{R}}\circ\mathcal{N}\right)(\sigma)=\sigma$, and also if $\rho$ and $(\mathcal{P}^{\mathrm{R}}\circ\mathcal{N})(\rho)$ are close then $-\log F\left(\rho,(\mathcal{P}^{\mathrm{R}}\circ\mathcal{N})(\rho)\right)$ will be a small number. So, the above refinement says that it is possible to perform a recovery operation in which one can perfectly recover one state and approximately recover the other states if the change in the quantum relative entropy between the two quantum states before and after the action of the quantum channel is small Wil (15); JRS+ (18).
@@
Let $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a superchannel in the class $\mathsf{SCT}$ (see Sec. II) and $\mathcal{N}_{A\rightarrow A^{\prime}}$, and $\mathcal{M}_{A\rightarrow A^{\prime}}$ be two quantum channels. Let $\Psi_{\widetilde{A}A}\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$ and $\Phi_{\widetilde{C}C}\in\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$ be the pure states that realize the supremum in the relative entropy functional $D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]$ and $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$, respectively. Then we have